5x+2=-x+14
6x-3x+2=-10
2(x+2)=3(x-5)
3(m-2)=-2(m-7)
6x+4x-2=15

Answer by checkley79(3341)

5x+2=-x+14
5x+x=14-2
6x=12
x=12/6
x=2 ans
--------------
6x-3x+2=-10
3x=-10-2
3x=-12
x=-12/3
x=-4 ans.
-------------
2(x+2)=3(x-5)
2x+4=3x-15
2x-3x=-15-4
-x=-19
x=19 ans.
----------------
I'll leave the rest for you to practice with.
3(m-2)=-2(m-7)
6x+4x-2=15